# Holographic Integrated Data and Energy Transfer

### v1

@@ III-A Near-filed Focus for Single EU
In the traditional near-field channel, the radius of the transmitting aperture is far less than the distance between the transmitter and the receiver. By applying Taylor series expansion $\sqrt{1+x}=1+x/2+\mathcal{O}(x^{2})$ in the traditional near-field channel [30], the near-field focusing is realized in the Fresnel region which ranges from the radius of $\frac{1}{2}\sqrt{\frac{D^{3}}{\lambda}}$ to $2\frac{D^{2}}{\lambda}$. Different from the traditional channel, the Fresnel region in the EM channel is extended below the lower bound $\frac{1}{2}\sqrt{\frac{D^{3}}{\lambda}}$, where the EUs are more likely to be located. In order to highlight the difference between the traditional channel and the EM channel as well as analyze the near-filed focusing effect in the H-IDET system, we firstly investigate the maximum energy harvesting amount in the near-field region in this sub-section. In other words, we aim to maximize the actual harvested energy after the non-linear EH model, which is the same as maximizing the input energy harvesting power at the receiver.
@@
According to Cauchy-Schwarz inequality $\left|\bm{\psi}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}(\mathbf{s})\bm{\theta}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}\leq\left|\int_{\mathcal{S}_{\mathrm{T}}}\left\|\bm{\theta}(\mathbf{s})\right\|^{2}\mathrm{d}\mathbf{s}\right|\left|\int_{\mathcal{S}_{\mathrm{T}}}\left\|\bm{\psi}^{H}\mathbf{G}(\mathbf{s})\right\|^{2}\mathrm{d}\mathbf{s}\right|$, when $\bm{\psi}$ is fixed, $\mathrm{P_{EH}}$ is maximized when
@@ III-B2 Problem Transformation
$=$ $\,{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}})}{\bm{\theta}_{k}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}-1}\right|^{2}}+\sum\limits_{j=1,j\neq k}^{K+L}{{{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}}){{\bm{\theta}}_{j}}\left({{\mathbf{s}}}\right)}{\mathrm{d}}{\mathbf{s}}}}\right|}^{2}}}\!+\!{\sigma^{2}}{\left\|{{{{\bm{\psi}}_{k}}}}\right\|^{2}}.$ (16)
@@
[…]
@@
$\bm{\psi}_{k}=\left(\sum_{j=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}+\sigma^{2}\mathbf{I}_{3}\right)^{-1}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s},$ (18)
which is also the optimum $\bm{\psi}_{k}$ for (P3). By substituting (18) into (III-B2), the corresponding MSE is obtained as
$M_{k}^{\mathrm{min}}=1-\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}\left(\sum_{j=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}\!\!\!+\sigma^{2}\mathbf{I}_{3}\right)^{-1}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}.$ (19)
@@
By substituting (21), (19) and (III-B2) into (P3) and utilizing $\mathrm{det}|\mathbf{I}+\mathbf{AB}|=\mathrm{det}|\mathbf{I}+\mathbf{BA}|$ and the matrix inversion lemma, i.e., $(\mathbf{A}+\mathbf{BCD})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\mathbf{B}(\mathbf{C}^{-1}+\mathbf{D}\mathbf{A}^{-1}\mathbf{B})^{-1}\mathbf{D}\mathbf{A}^{-1}$, the object of (P3) is derived as
@@ III-B3 Holographic Beamforming Design
The block coordinate descent (BCD) algorithm is conceived to solve problem (P3), where the variables are separated and optimized iteratively by fixing the other one as the constant. Specifically, given the beamforming set $\bm{\Theta}(\mathbf{s})$, $\bm{\rho}$ and $\bm{\Psi}$ are already obtained by (18), (III-B2) and (21) respectively in the proof of Lemma 1. In the context, we aim to optimize $\bm{\Theta}(\mathbf{s})$ by fixing $\bm{\rho}$ and $\bm{\Psi}$, while the problem (P3) is then re-formulated as
(P4): $\min\limits_{\bm{\Theta(\mathbf{s})}}\sum\limits_{k=1}^{K}\rho_{k}\left(\sum_{j=1}^{K+L}\left|\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}-2\mathfrak{R}\left\{\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}\right)$ (23)
@@
where ${\Upsilon_{\mathbf{n}}}\left({{\mathbf{s}}}\right)=\frac{1}{\sqrt{A_{\mathrm{T}}}}{e^{2\pi{\mathrm{j}}\left({{\frac{n_{x}}{L_{x}}\left({s_{x}-\frac{L_{x}}{2}}\right)+\frac{n_{y}}{L_{y}}{\left({s_{y}-\frac{L_{y}}{2}}\right)}+\frac{n_{z}}{L_{z}}{\left({s_{z}-\frac{L_{z}}{2}}\right)}}}\right)}}\in\mathbb{C}$ is the Fourier orthogonal base function, and ${{\mathbf{w}}_{k,{\mathbf{n}}}}=\frac{1}{\sqrt{A_{\mathrm{T}}}}\int_{{\mathbf{s}}\in{\mathcal{S}}_{\mathrm{T}}}{\bm{\theta}_{k}({\mathbf{s}}){\Upsilon^{\ast}_{\mathbf{n}}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}\in\mathbb{C}^{3\times 1}$ is the projection length in the wavenumber domain. For the simplicity, we define $\mathbf{n}:=(n_{x},n_{y},n_{z})\in\mathbb{Z}^{3\times 1}$, $\mathbf{N}:=(N_{x},N_{y},N_{z})\in\mathbb{Z}^{3\times 1}$, $\sum_{\mathbf{n}}^{\mathbf{N}}:=\sum_{n_{x}=-N_{x}}^{N_{x}}\sum_{n_{y}=-N_{y}}^{N_{y}}\sum_{n_{z}=-N_{z}}^{N_{z}}$. The total number of Fourier bases is $N_{F}=(2N_{x}+1)(2N_{y}+1)(2N_{z}+1)$. Besides, $L_{x}$, $L_{y}$ and $L_{z}$ denote the maximum projection lengths of $\mathcal{S}_{\mathrm{T}}$ on 3D cartesian coordinate system, respectively. Then, in the electric field and power constraint, Fourier transform can be applied as [23]:
@@
where $\mathbf{h}_{k}$ and $\mathbf{w}_{k}$ are the vectorized sets of $\mathbf{h}_{k,\mathbf{n}}$ and $\mathbf{w}_{k,\mathbf{n}}$ for all $\mathbf{n}\in\{\{-N_{x},\cdots,N_{x}\},\\ \{-N_{y},\cdots,N_{y}\},\{-N_{z},\cdots,N_{z}\}\}$ respectively, i.e., $\mathbf{h}_{k}=\left[\mathbf{h}_{k,\mathbf{1}};\dots;\mathbf{h}_{k,\mathbf{N}}\right]$ and $\mathbf{w}_{k}=\left[\mathbf{w}_{k,\mathbf{1}};\dots;\mathbf{w}_{k,\mathbf{N}}\right]$. $\mathbf{h}_{k,\mathbf{n}}:=\bm{\Omega}_{k,{\mathbf{n}}}^{H}\bm{\psi}_{k}$ and $\bm{w}:=\{\mathbf{w}_{k}\}_{k=1}^{K+L}$.
The problem (P5) is a quadratically constrained quadratic programming (QCQP). However, the energy harvesting constraints (28b) are still non-convex, so that the successive convex approximation (SCA) method is then employed for obtaining the solution. Given the beamfroming vector $\mathbf{\bar{w}}_{j}$, the first-order Taylor expansion is expressed as

### v2

@@ III-A Near-filed Focus for Single EU
In the traditional near-field channel, the radius of the transmitting aperture is far less than the distance between the transmitter and the receiver. By applying Taylor series expansion $\sqrt{1+x}=1+x/2+\mathcal{O}(x^{2})$ in the traditional near-field channel [38], the near-field focusing is realized in the Fresnel region which ranges from the radius of $\frac{1}{2}\sqrt{\frac{D^{3}}{\lambda}}$ to $2\frac{D^{2}}{\lambda}$ [39]. Different from the traditional channel, the Fresnel region in the EM channel is extended below the lower bound $\frac{1}{2}\sqrt{\frac{D^{3}}{\lambda}}$, where the EUs are more likely to be located. In order to highlight the difference between the traditional channel and the EM channel as well as analyze the near-filed focusing effect in the H-IDET system, we firstly investigate the maximum energy harvesting amount in the near-field region in this sub-section. In other words, we aim to maximize the actual harvested energy after the non-linear EH model, which is the same as maximizing the input energy harvesting power at the receiver, since $\Xi(P_{l})$ in Eq. (10) monotonically increases with respect to $P_{l}$.
@@
According to the Cauchy-Schwarz inequality :
$\left|\bm{\psi}^{H}\!\!\int_{\mathcal{S}_{\mathrm{T}}}\!\!\mathbf{G}(\mathbf{s})\bm{\theta}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}\!\!\leq\left|\int_{\mathcal{S}_{\mathrm{T}}}\!\!\left\|\bm{\theta}(\mathbf{s})\right\|^{2}\mathrm{d}\mathbf{s}\right|\left|\int_{\mathcal{S}_{\mathrm{T}}}\!\!\left\|\bm{\psi}^{H}\mathbf{G}(\mathbf{s})\right\|^{2}\mathrm{d}\mathbf{s}\right|,$ (13)
when $\bm{\psi}$ is fixed, $\mathrm{P_{EH}}$ is maximized when
@@
In Section III-A, the obtained closed-form beamforming can achieve single-point near-field focusing for WET. In Section III-B, the proposed BCD-based beamforming scheme will achieve multi-point near-field focusing simultaneously for the H-IDET system.
@@ III-B2 Problem Transformation
$=$ $\,{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}})}{\bm{\theta}_{k}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}-1}\right|^{2}}$
$+$ $\sum\limits_{j=1,j\neq k}^{K+L}{{{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}}){{\bm{\theta}}_{j}}\left({{\mathbf{s}}}\right)}{\mathrm{d}}{\mathbf{s}}}}\right|}^{2}}}\!+\!{\sigma^{2}}{\left\|{{{{\bm{\psi}}_{k}}}}\right\|^{2}}.$ (17)
@@
where $P_{0}^{{}^{\prime}}=\frac{A_{\text{R}}\cos(\phi)}{Z}\Xi^{-1}(P_{0})$ and $\Xi^{-1}$ is inverse function of $\Xi$ in Eq. (10).
@@
$\bm{\psi}_{k}$ $=\left(\sum_{j=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}+\sigma^{2}\mathbf{I}_{3}\right)^{-1}$
×$\times\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s},$ (19)
which is also the optimum $\bm{\psi}_{k}$ for (P3). By substituting (III-B2) into (III-B2), the corresponding MSE is obtained as
$M_{k}^{\mathrm{min}}$ $=-\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}$
×$\times\left(\sum_{j=1}^{K+L}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\left(\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right)^{H}\!\!\!+\sigma^{2}\mathbf{I}_{3}\right)^{-1}$
×$\times\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}+1.$ (20)
@@
By substituting (22), (III-B2) and (III-B2) into (P3) and utilizing $\mathrm{det}|\mathbf{I}+\mathbf{AB}|=\mathrm{det}|\mathbf{I}+\mathbf{BA}|$ and the matrix inversion lemma, i.e., $(\mathbf{A}+\mathbf{BCD})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\mathbf{B}(\mathbf{C}^{-1}+\mathbf{D}\mathbf{A}^{-1}\mathbf{B})^{-1}\mathbf{D}\mathbf{A}^{-1}$, the object of (P3) is derived as
@@ III-B3 Holographic Beamforming Design
The block coordinate descent (BCD) algorithm is conceived to solve problem (P3), where the variables are separated and optimized iteratively by fixing the other one as the constant. By introducing the auxiliary variables $\rho$ in (P3) and based on BCD algorithm, the subproblem of (P3) without $\log$ function can be optimized. Specifically, given the beamforming set $\bm{\Theta}(\mathbf{s})$, $\bm{\rho}$ and $\bm{\Psi}$ are already obtained by (III-B2), (III-B2) and (22) respectively in the proof of Lemma 1. In the context, we aim to optimize $\bm{\Theta}(\mathbf{s})$ by fixing $\bm{\rho}$ and $\bm{\Psi}$. By substituting the MSE $M_{k}$ in (III-B2) into (18) and discarding the constant terms, we obtain
$\sum\limits_{k=1}^{K}\rho_{k}\left(\,{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}})}{\bm{\theta}_{k}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}-1}\right|^{2}}\!\!+\!\!\sum\limits_{j=1,j\neq k}^{K+L}{{{\left|{{\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}}){{\bm{\theta}}_{j}}\left({{\mathbf{s}}}\right)}{\mathrm{d}}{\mathbf{s}}}}\right|}^{2}}}\right)$
$=$ $\sum\limits_{k=1}^{K}\rho_{k}\left(\sum_{j=1}^{K+L}\left|\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\!\!\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}\!\!\!-\!2\mathfrak{R}\left\{\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\!\!\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}\right)+\sum\limits_{k=1}^{K}\rho_{k}$ (24)
where the real term $2\mathfrak{R}\left\{\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\!\!\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}$ is obtained form ${\bm{\psi}}_{k}^{H}{\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}})}{\bm{\theta}_{k}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}+\left({\int_{{S_{\mathrm{T}}}}{{{\mathbf{G}}_{k}}({\mathbf{s}})}{\bm{\theta}_{k}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}\right)^{H}{\bm{\psi}}_{k}$. By discarding the constant terms in (III-B3), the problem (P3) is then re-formulated as
(P4): $\min\limits_{\bm{\Theta(\mathbf{s})}}\sum\limits_{k=1}^{K}\rho_{k}\Bigg{(}\sum_{j=1}^{K+L}\left|\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right|^{2}$
$\ \ \ \ \ -2\mathfrak{R}\left\{\bm{\psi}_{k}^{H}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{k}(\mathbf{s})\bm{\theta}_{k}(\mathbf{s})\mathrm{d}\mathbf{s}\right\}\Bigg{)}$ (25)
@@
where ${\Upsilon_{\mathbf{n}}}\left({{\mathbf{s}}}\right)=\frac{1}{\sqrt{A_{\mathrm{T}}}}{e^{2\pi{\mathrm{j}}\left({{\frac{n_{x}}{L_{x}}\left({s_{x}-\frac{L_{x}}{2}}\right)+\frac{n_{y}}{L_{y}}{\left({s_{y}-\frac{L_{y}}{2}}\right)}+\frac{n_{z}}{L_{z}}{\left({s_{z}-\frac{L_{z}}{2}}\right)}}}\right)}}\in\mathbb{C}$  [^3] The orthogonal basis has a similar form mathematically to the DFT beam. However, the beamforming $\bm{\theta}(\mathbf{s})$ is expanded for each point on the continuous aperture and the number of bases $N_{F}$ is much smaller than the number of points. Moreover, the basis functions only provide functional degrees of freedom for the further optimization of beamforming functions $\bm{\theta}(\mathbf{s})$, which have no physical significance. is the Fourier orthogonal base function, and ${{\mathbf{w}}_{k,{\mathbf{n}}}}=\frac{1}{\sqrt{A_{\mathrm{T}}}}\int_{{\mathbf{s}}\in{\mathcal{S}}_{\mathrm{T}}}{\bm{\theta}_{k}({\mathbf{s}}){\Upsilon^{\ast}_{\mathbf{n}}}\left({{\mathbf{s}}}\right){\mathrm{d}}{\mathbf{s}}}\in\mathbb{C}^{3\times 1}$ is the projection length in the wavenumber domain. For the simplicity, we define $\mathbf{n}:=(n_{x},n_{y},n_{z})\in\mathbb{Z}^{3\times 1}$, $\mathbf{N}:=(N_{x},N_{y},N_{z})\in\mathbb{Z}^{3\times 1}$, $\sum_{\mathbf{n}}^{\mathbf{N}}:=\sum_{n_{x}=-N_{x}}^{N_{x}}\sum_{n_{y}=-N_{y}}^{N_{y}}\sum_{n_{z}=-N_{z}}^{N_{z}}$. The total number of Fourier bases is $N_{F}=(2N_{x}+1)(2N_{y}+1)(2N_{z}+1)$. Besides, $L_{x}$, $L_{y}$ and $L_{z}$ denote the maximum projection lengths of $\mathcal{S}_{\mathrm{T}}$ on 3D cartesian coordinate system, respectively. Then, in the electric field and power constraint, Fourier transform can be applied as [27]:
@@
 [^4] The conventional DFT beam can only be used in the far field because the channel response vector linear with respect to the antenna index is used. However, the channel characterization in wavenumber space is applicable in both the far-field and near-field.
@@
Remark 2
Thanks to the inherent physical properties of electromagnetic channel function $\mathbf{G}_{k}(\mathbf{s})$, some recent works have revealed that, the value of channel gain $\left\|\bm{\Omega}_{k,{\mathbf{n}}}\right\|^{2}$ in the wavenumber domain, is high in the low-wavenumber band and is negligible in the high-wavenumber band in most cases [26, 27]. Therefore, the functions $\left\{\bm{\theta}_{k}(\mathbf{s})\right\}_{k=1}^{K+L}$ and $\left\{\mathbf{G}_{k}(\mathbf{s})\right\}_{k=1}^{K+L}$ can be approximated by finite low-wavenumber and high-power items. Specifically, when $N_{x}=\lceil\frac{L_{x}}{\lambda}\rceil$, $N_{y}=\lceil\frac{L_{y}}{\lambda}\rceil$ and $N_{z}=\lceil\frac{L_{z}}{\lambda}\rceil$, i.e., the number of finite terms $N_{F}$ is not less than $(2\lceil\frac{L_{x}}{\lambda}\rceil+1)(2\lceil\frac{L_{y}}{\lambda}\rceil+1)(2\lceil\frac{L_{z}}{\lambda}\rceil+1)$, there are sufficient functional degrees of freedom for approximation [26, 27]. Thus, the performance loss can be artificially controlled by presetting an acceptable number of finite terms $N_{F}$.
@@
where $\mathbf{h}_{k}$ and $\mathbf{w}_{k}$ are the vectorized sets of $\mathbf{h}_{k,\mathbf{n}}$ and $\mathbf{w}_{k,\mathbf{n}}$ for all $\mathbf{n}\in\{\{-N_{x},\cdots,N_{x}\},\{-N_{y},\cdots,N_{y}\},\{-N_{z},\cdots,N_{z}\}\}$ respectively, i.e., $\mathbf{h}_{k}=\left[\mathbf{h}_{k,\mathbf{1}};\dots;\mathbf{h}_{k,\mathbf{N}}\right]$ and $\mathbf{w}_{k}=\left[\mathbf{w}_{k,\mathbf{1}};\dots;\mathbf{w}_{k,\mathbf{N}}\right]$. $\mathbf{h}_{k,\mathbf{n}}:=\bm{\Omega}_{k,{\mathbf{n}}}^{H}\bm{\psi}_{k}$ and $\bm{w}:=\{\mathbf{w}_{k}\}_{k=1}^{K+L}$.
Remark 3
By using the Fourier transform with finite number of bases, the functional programming (P4) is transformed into the classic quadratically constrained quadratic programming (QCQP) (P5). Therefore, the classic schemes for solving the QCQP can be adopted to solve it. However, there is a performance gap between (P4) and (P5), which is dependent with the number $N_{F}$ of orthogonal bases. Since $N_{F}$ represents the accuracy of approximation in Eq. (26), the gap can be narrowed as $N_{F}$ increases, which is also shown in Fig. (6). Besides, a lager $N_{F}$ also results in a higher computational complexity, which is showed in the complexity analysis in Section III-(5).
Furthermore, different from the physical channel in conventional problems, $\mathbf{h}_{k}$ in (P5) is the equivalent counterpart of the physical electromagnetic channel after the Fourier transform and polarization combining. Moreover, the variable $\bm{w}$ in (P5) is the projection length of $\bm{\theta}(\mathbf{s})$ on different orthogonal bases, while the variables in conventional problems are usually matrix of digital precoder and phase shifters.
Besides, conventional MIMO systems consist of discrete antenna arrays with half wavelength spacing and is fed in parallel. Holographic MIMO systems use metasurface antennas and break through the limit of half wavelength spacing. The near-continuous aperture enables powerful electromagnetic manipulation capabilities, which can further approach the performance limit. The practical hardware of holographic MIMO mostly uses series feeding, which has higher energy efficiency.
The problem (P5) is a QCQP. However, the energy harvesting constraints (30b) are still non-convex, so that the successive convex approximation (SCA) method is then employed for obtaining the solution. Given the beamfroming vector $\mathbf{\bar{w}}_{j}$, the first-order Taylor expansion is expressed as